Home   >  Maths quizzes   >  Grade 5 Quizzes   >  Probability trees

Probability trees Quiz

Grade 5 • 5W • 10 Questions (39 Marks)

Score: 0 / 10
Probability trees — Notes & Worked Examples Show / Hide

Explanation:

On a probability tree diagram, each pair of branches from the same point must add up to 1. Multiply along the branches for AND (P(A and B) = P(A) × P(B)), and add the end probabilities together when more than one path satisfies the question (OR).

Check whether items are picked with replacement (second-pick probabilities stay the same) or without replacement (reduce total denominator by 1 on the second branches, and reduce the numerator by 1 if that colour was already picked).

Example 1: The probability Sam passes his theory test is 0.8 and practical test is 0.6. Find the probability he passes only one test.

Solution:

P(P, F) = 0.8 × 0.4 = 0.32; P(F, P) = 0.2 × 0.6 = 0.12
Answer: 0.32 + 0.12 = 0.44
Theory Practical 0.8 0.2 P F 0.6 0.4 0.6 0.4 P (0.48) F → 0.32 ✓ P → 0.12 ✓ F (0.08)

Example 2: A bag has 5 red and 3 blue counters. Two are picked without replacement. Find P(both same colour).

Solution:

P(Red, Red) = 58 × 47 = 2056
P(Blue, Blue) = 38 × 27 = 656
Total = 2056 + 656 = 2656 = 1328
Answer: 1328
Question 1
[2 marks]
RedBlueRedBlueRedBlue0.30.70.30.70.30.7
Work out
A bag contains Red and Blue counters.
A counter is picked, replaced, and another is picked.
Find the probability of picking exactly 1 Red counter.
Worked Solution & Answer
Probability = successful outcomestotal possible outcomes (all mutually exclusive outcomes add up to 1).
0.42
Answer: 0.42
Question 2
[2 marks]
A biased coin is flipped twice.
The probability of Heads is 0.3.
HTHTHT
(a) Complete the probability tree. [1 mark]
(b) What is the probability of getting two Heads? [1 mark]
Worked Solution & Answer
(a)
Branches: 0.3, 0.7
Answer (a): Branches: 0.3, 0.7
(b)
0.09
Answer (b): 0.09
Question 3 AQA Jun 2025 2F Q19
[3 marks]
Two spinners each have only red sections and green sections.
Spinner 1: P(red) = 38
Spinner 2: P(red) = 15
(a) Complete the tree diagram. [2 marks]
Spinner 1Spinner 2RedGreen⅜RedGreen⅕RedGreen⅕
(b) Both spinners are spun.
Work out the probability that both spinners land on red.
[1 mark]
Answer
Worked Solution & Answer
(a)
Spinner 1 green = 1 – 3/8 = 5/8. Spinner 2 green = 1 – 1/5 = 4/5 M1 A1
Spinner 1Spinner 2RedGreen⅜⅝RedGreen⅕⅘RedGreen⅕⅘
Answer (a): Spinner 1 Green = 5/8; Spinner 2 Green = 4/5 (in both positions)
(b)
P(Red and Red) = 3/8 × 1/5 = 3/40 = 0.075 B1
Answer (b): Answer 3/40
Question 4 Eduqas Jun 2024 2F Q19 / 2H Q4
[4 marks]
For her lunch, Gina sometimes eats a salad and sometimes drinks a coffee.
The probability she eats a salad for lunch on a given day is 0.4.
The probability she drinks a coffee on a given day is 0.25.
Eating a salad for lunch and drinking a coffee are independent.
(a) Complete the following tree diagram. [2 marks]
0·4……...Gina eats a saladfor lunchGina does not eata salad for lunch0·25……...Gina drinks a coffeeGina does not drink a coffee……...……...Gina drinks a coffeeGina does not drink a coffee
(b) Calculate the probability that, on a given day, Gina does not eat a salad for lunch but does drink a coffee. [2 marks]
Worked Solution & Answer
(a)
P(does not eat salad) = 1 - 0.4 = 0.6. M1
Since drinking coffee is independent: P(drinks coffee) = 0.25, P(does not drink coffee) = 1 - 0.25 = 0.75 on both salad and no-salad branches. A1
0·40·6Gina eats a saladfor lunchGina does not eata salad for lunch0·250·75Gina drinks a coffeeGina does not drink a coffee0·250·75Gina drinks a coffeeGina does not drink a coffee
Answer (a): P(not salad) = 0.6; P(coffee) = 0.25 and P(not coffee) = 0.75 on both branch sets.
(b)
P(No salad AND Drinks coffee) = 0.6 × 0.25 = 0.15 M1 A1
Answer (b): 0.15
Question 5
[4 marks]
Tim has two biased coins, coin **A** and coin **B**.
He is going to throw both coins.
The probability that coin **A** will land on heads is 0.6
The probability that coin **B** will land on heads is 0.55
(a) Complete the probability tree diagram. [2 marks]
coin Acoin Bheadsnot headsheadsnot headsheadsnot heads0.60.55
(b) Work out the probability that both coins land on heads. [2 marks]
Worked Solution & Answer
(a)
1 – 0.6 = 0.4 M1
1 – 0.55 = 0.45 A1
coin Acoin Bheadsnot headsheadsnot headsheadsnot heads0.60.550.40.450.550.45
Answer (a): Coin A: 0.4; Coin B: 0.45, 0.55, 0.45
(b)
0.6 × 0.55 = 0.33 M1 A1
Answer (b): 0.33
Question 6 IGCSE Nov 2024 1H Q13
[4 marks]
The probability that Thomas gets to work early on Saturday is 0.7
If Thomas gets to work early on Saturday, the probability that he will get to work early on Sunday is 0.9
If Thomas does not get to work early on Saturday, the probability that he will get to work early on Sunday is 0.6
(a) Use this information to complete the probability tree diagram. [2 marks]
SaturdaySunday0.7...............earlynotearly..............................earlynotearly..............................earlynotearly
(b) Work out the probability that Thomas gets to work early on both Saturday and Sunday. [2 marks]
Worked Solution & Answer
(a)
Saturday not early = 0.3 M1
Sunday branches = 0.9, 0.1 and 0.6, 0.4 A1
SaturdaySunday0.70.3earlynotearly0.90.1earlynotearly0.60.4earlynotearly
Answer (a): Saturday not early = 0.3; Sunday branches = 0.9, 0.1 and 0.6, 0.4
(b)
0.7 × 0.9 = 0.63 M1 A1
Answer (b): 0.63
Question 7
[5 marks]
Freya and Ria each do an exam.
The probability that Freya will pass the exam is 0.8
The probability that Ria will pass the exam is 0.85
(a) Complete the probability tree diagram. [2 marks]
FreyaRiapassfailpassfailpassfail0.80.85
(b) Work out the probability that only one of Freya or Ria will pass the exam. [3 marks]
Worked Solution & Answer
(a)
Freya fail = 1 – 0.8 = 0.2. Ria fail = 1 – 0.85 = 0.15 M1 A1
FreyaRiapassfailpassfailpassfail0.80.850.20.150.850.15
Answer (a): Tree completed (Freya fail = 0.2, Ria pass = 0.85, Ria fail = 0.15)
(b)
P(Freya pass, Ria fail) = 0.8 × 0.15 = 0.12 M1
P(Freya fail, Ria pass) = 0.2 × 0.85 = 0.17 M1
P(only one) = 0.12 + 0.17 = 0.29 A1
Answer (b): 0.29
Question 8 AQA Jun 2024 3H Q12
[5 marks]
Cards are either gold or not gold.
P(gold) = 0.05
Harim chooses a card at random and replaces it.
He then chooses a second card.
(a) Complete the tree diagram. [2 marks]
First cardSecond card0.05GoldNot gold0.05GoldNot goldGoldNot gold
(b) What is the probability that **at least one** of Harim’s cards is gold? [3 marks]
Answer
Worked Solution & Answer
(a)
P(Not gold) = 1 – 0.05 = 0.95 M1 A1
First cardSecond card0.05GoldNot gold0.05GoldNot goldGoldNot gold0.950.950.050.95
Answer (a): First card Not gold = 0.95; Second card branches: 0.95, and 0.05 / 0.95
(b)
P(both not gold) = 0.95 × 0.95 = 0.9025 M2
P(at least one gold) = 1 – 0.9025 = 0.0975 A1
Answer (b): Answer 0.0975
Question 9
[5 marks]
Mary travels to work by train every day.
The probability that her train will be late on any day is 0.15
(a) Complete the probability tree diagram for Thursday and Friday. [2 marks]
ThursdayFridaylatenot late0.15latenot latelatenot late
(b) Work out the probability that her train will be late on at least one of these two days. [3 marks]
Worked Solution & Answer
(a)
0.85 on Thursday not late M1
0.15 and 0.85 on both Friday pairs A1
ThursdayFridaylatenot late0.150.85latenot late0.150.85latenot late0.150.85
Answer (a): 0.85 on Thursday not late; 0.15 and 0.85 on both Friday pairs
(b)
1 – P(not late, not late) = 1 – (0.85 × 0.85) = 1 – 0.7225 = 0.2775 M2 A1
Answer (b): 0.2775
Question 10 OCR Jun 2025 2F Q25
[5 marks]
Casey plays two games of football.
In each game, the probability that Casey scores a goal is 0.3.
(a) Complete the tree diagram. [2 marks]
First gameSecond gameScoresDoes notscoreScoresDoes notscoreScoresDoes notscore
(b) Work out the probability that Casey scores in only one of the games. [3 marks]
Worked Solution & Answer
(a)
1 − 0.3 = 0.7. Label each Scores branch 0.3 and each Does not score branch 0.7. M1 A1
First gameSecond gameScoresDoes notscoreScoresDoes notscoreScoresDoes notscore0.30.70.30.70.30.7
Answer (a): 0.3 on all Scores branches and 0.7 on all Does not score branches
(b)
P(Scores then Does not score) = 0.3 × 0.7 = 0.21 M1
P(Does not score then Scores) = 0.7 × 0.3 = 0.21 M1
Total probability = 0.21 + 0.21 = 0.42. A1
Answer (b): 0.42
Score: 0 / 10 (0 / 39 marks)
Probability trees Worksheets Next: Vectors →